OpenStudy (hoybaby):

how do you find f(g(n)) if f(n)=3n-1 and g(n) = n^2 + 5. Please help i am totally confused

OpenStudy (anonymous):

plug n^2 + 5 in f(n)

OpenStudy (anonymous):

===> f(g(n))= 3(n^2 + 5)-1

jimthompson5910 (jim_thompson5910):

Start with f(n) f(n)=3n-1 Now replace EVERY copy of 'n' with g(n) like so f(g(n))=3*g(n)-1 Now, because g(n) = n^2+5, this means that we can replace the "g(n)" on the right side with n^2+5 to get f(g(n)) = 3(n^2+5)-1 Finally, expand and simplify as must as possible f(g(n)) = 3(n^2+5)-1 f(g(n)) = 3(n^2)+3(5)-1 f(g(n)) = 3n^2+15-1 f(g(n)) = 3n^2+14